Integral (x+2)dx/(2x^2 +4x+3)^2 plz solve I m very confused

Grade:

1 Answers

Khimraj
6 years ago
I = ∫(x+2)dx/(2x2 +4x+3)2
put 2x2 + 4x +3 = t
then (4x + 4)dx = dt
(2x + 2)dx = dt/2
Then I = ½∫dt/t2 = ½(-1/t) + c
So I = -½(1/(2x2 + 4x +3)) +c.
Hope it clears. If you like answer then please approve it.